# Unification of the Gauge Theories

### v2

@@ VI Discussion
Here, we have used a quaternion mechanism to form new vectors. We have provided $q$-vectors as physical quantities—such as events in the manifold of spacetime—that correspond to the position of the Lorentz indices and are coupled to the generators of quaternion algebra. By introducing the operator $\hat{\Gamma}_{\mu}$ and the mapping "$\mathsf{ast}"$ ($\ast$-product), we then succeeded in obtaining the theory of gauge fields in a general spacetime. We introduce the new antisymmetric tensor $n_{k}^{\ ij}$ and use the equations of motion of gauge theory to obtain the Bianchi identities. We have also shown that the Bianchi identities are physical relations, not mathematical structures. As we proceeded, we replaced the gauge fields with the Christoffel coefficients and followed the improved quaternion mechanism to access the theory of gravity. Our calculations indicate that with this substitution and additional degrees of freedom for Riemann’s curvature tensor, the theory of gravitation remains valid, and the Einstein constant term is essential for the equations of motion. The results of gravity are much higher than those of electrodynamic theory, and the reason for this difference is Riemann’s curvature four-component tensor. As for the Bianchi identities in gravity, we have concluded that the identities are part of the equations of motion. One of the most significant outcomes of deriving the equation of gravity is the formalization of the constant term, referred to as Einstein’s constant term, in the Einstein equation. We have also established that the constant term is exclusively geometric. Having decomposed the four-component Riemann’s curvature tensor into one with two component subsections, another result of this work is to emphasize the independent behavior of the two-component subsections of the curvature tensor. The result of the present work is a commentary on parisi ; faisal and many other references and matches the physical data with the solutions of the Einstein equation, with many corrections to the Einstein theory of gravity using the unprincipled modification approach to the Einstein equation (modifications that do not apply to equation 34). By analyzing the components of Riemann’s curvature tensor, we can develop a new approach to quantum gravity. The conclusion of this article is that any manipulation of Einstein’s equation is incorrect and must comply with the condition ”the covariant derivative equals zero.”

### v3

@@ V Discussion
Here, we have used a quaternion mechanism to form new vectors. We have provided $q$-vectors as physical quantities, events in the manifold of spacetime that correspond to the position of the Lorentz indices, which have coupling with the generators of quaternion algebra. Employing operator $\hat{\Gamma}_{\mu}$ and mapping $\mathsf{ast}$ ($\ast$-product), we successfully developed the theory of gauge fields in a general spacetime. We introduce the new antisymmetric tensor $n_{k}^{\ ij}$ and use the equations of motion of gauge theory to obtain the Bianchi identities. We have demonstrated that the Bianchi identities represent physical equations of motion rather than merely mathematical relationships. As we proceeded, we replaced the gauge fields with the Christoffel coefficients and followed the improved quaternion mechanism to access the theory of gravity. Our calculations indicate that with this substitution and assuming more degrees of freedom for Riemann’s curvature tensor, the theory of gravitation remains valid, and Einstein’s constant term for the equations of motion can be formally defined. The results of the theory of gravity are much higher than those of electrodynamic theory, and the reason is Riemann’s curvature four-component tensor. As for the Bianchi identities in gravity, we have concluded that the identities are part of the equations of motion. One of the most significant outcomes of deriving the equation of gravity is the formalization of the constant term, referred to as Einstein’s constant term, in the Einstein equation. We have also established that the constant term is exclusively geometric. Having decomposed the four-component Riemann’s curvature tensor into one with two component subsections, another result of this work is to emphasize the independent behavior of the two-component subsections of the curvature tensor. The result of the present work is a commentary on parisi ; faisal . Many other references fit the physical data with the solutions of Einstein’s equation, along with introducing many modifications to Einstein’s theory of gravity and using non-principled correction approaches to Einstein’s equation (corrections that are not applied to the equation 34). By analyzing the components of Riemann’s curvature tensor, we develop a quaternion approach and we can suggest a new way to quantum gravity, with additional equations of motion. This article concludes that any manipulation of Einstein’s equation is wrong except to meet the condition that the covariant derivative is zero.
VI Acknowledgments